if the nine-digit integer 3365097 ab is divisible by 72. find the possible ordered pairs of digits (a,b)
Answer by ankor@dixie-net.com(22740)

if the nine-digit integer 3365097 ab is divisible by 72.
find the possible ordered pairs of digits (a,b)
:
try a=9, b=9
Using a calc 336509799/72 = 4673747.208
:
.208 * 72 ~ 15
:
Subtract 15 from 99 = 84
a=8, b=4
336509784/72 = 4673747
:
Subtract 72:
a=1, b=2
336509712/72 = 4673746
;
Possible a, b values
8, 4 and 1, 2